### v1

# Dispersion-Mediated Space-Time States

Abstract
Space-time varying media enable unprecedented control over electromagnetic waves, yet most existing studies assume idealized, nondispersive materials and thus fail to capture the intrinsic frequency dispersion of realistic platforms. Here, we develop a general framework for dispersive space-time varying systems that rigorously identifies the physically allowed frequency transitions of waves scattered at uniformly moving interfaces. Unlike previous approaches, our method is valid for arbitrary dispersion profiles, including resonances, and does not rely on the commonly used frame hopping approach, allowing treatment of multiple-velocity and accelerated systems. Applying this framework to canonical Drude and Lorentz media, we uncover a family of dispersion-mediated space-time states that arise from the multiple frequency transitions permitted by material dispersion. These states extend beyond conventional nondispersive scattering, revealing qualitatively new regimes of space-time scattering regimes. Our results establish a rigorous foundation for the design of realistic space-time metamaterials, with immediate relevance to emerging experiments in epsilon-near-zero optical platforms and open pathways for dispersion-engineered wave manipulation.
I Introduction
Space-time varying media are advanced structures whose material properties, such as the refractive index, are modulated in both space and time [Cassedy1963_PUB, Cassedy1967_PUB, Caloz2019a_ST_Metamaterials_PUB, Caloz2019b_ST_Metamaterials_PUB], with the modulations typically produced by an external drive such as electronic, optical, acoustic or mechanical pumps [Saleh2019_BOOK, Rhodes1981_Signal_Proc_PUB, Shalaev2019_Active_Meta_PUB]. Compared to static media, space-time varying media support qualitatively different wave interactions and enable new physical effects [Caloz2022_GSTEMs_PUB], which in turn unlock a range of novel applications, including motion-induced photon cooling [Pendry2024_Air_Cond_Phot_USTEM_PUB], arbitrary pulse shaping [Bahrami2025_Pulse_Shap_ASTEM_PUB], generalized frequency chirping [DeKinder2025_Scat_Chirp_ASTEM], interface-based light amplification [Pendry2021_Gain_PUB, Luo2023_PUB], time reversal [Fink2016_TR_TEM_PUB], temporal beam splitting [Guerreiro2003_PUB] and Doppler pulse amplification [DeKinder2025_DoPA_USTEM_PUB].
While many of these applications have been explored in idealized, nondispersive media, physical materials inherently exhibit frequency-dependent dispersive responses that can strongly influence wave propagation, causing effects such as pulse distortion and frequency chirping [Saleh2019_BOOK]. Moreover, dispersion is particularly relevant in experiments, where it may dominate the system behavior over the idealized, nondispersive case. Recent works have begun addressing dispersive time-varying media [Fante1971_Transmission_TEM_PUB, Fante1973_Prop_EM_TEM_PUB, Pendry2022_Review_TEM_PUB, Engheta2025_Disp_Hyper_Media_TEM_PUB, Horsley2023_Eigenpulses_PUB, Engheta2021_Dispersion_PUB, Mirmoosa2022_Dipole_PUB, Joannopoulos2024_Opt_Prop_PUB, Maslov2021_Light_Scat_PUB, Maslov2021_Const_Rel_PUB, Maslov2024_Dispersion_PUB, Monticone2022_Dispersion_PUB], deriving new fundamental bounds, such as Kramers-Kronig relations [Kramers1927_KK_Relations_PUB, Kronig1926_KK_Relations_PUB] for time-varying media and exploring their impact on wave dynamics [Monticone2021_Spectral_Caus_PUB, Engheta2021_Kramers_Kronig_PUB].
Despite these advances, dispersive space-time varying systems have received comparatively little attention. Apart from a series of studies by the group of Agrawal [Agrawal2015_Temporal_Analog_USTEM_PUB, Agrawal2016_Spectral_Splitting_PUB, Agrawal2021_Refl_and_Trans_USTEM_PUB], most existing analyses are limited to purely temporal modulations. Moreover, the approaches in [Agrawal2015_Temporal_Analog_USTEM_PUB, Agrawal2016_Spectral_Splitting_PUB, Agrawal2021_Refl_and_Trans_USTEM_PUB] are limited by several assumptions: (1) polynomial Taylor approximation of dispersion relations, (2) slowly varying envelope approximation and (3) frame hopping to a comoving reference frame. These assumptions significantly limit their range of applicability, especially near resonances or in systems with multiple or nonuniform velocities. Meanwhile, experimental interest in space-time varying systems is rapidly growing, with $\epsilon$-near-zero (ENZ) materials emerging as a particularly promising platform in the optical regime [Shalaev2023_PTC_TEM_PUB, Boyd2020_Time_Refr_ENZ_TEM_PUB, Segev2023_Single_Cycle_TEM_PUB, Kinsey2025_ST_Knife_TEM_PUB]. Because ENZ materials are highly dispersive, accurately predicting the response of a moving modulation requires a systematic framework capable of handling fully dispersive space-time varying systems, including frequencies near resonance.
In this work, we provide such a framework by identifying the physically allowed frequency transitions of scattered waves at a uniformly moving interface separating two dispersive media, opening the pathway for generalized dispersion engineering. We show that dispersion can create additional frequency-transition branches that have no counterpart in the nondispersive limit, giving rise to new “dispersion-mediated space-time states”, depending on the dispersive properties of the media, the interface velocity and the incident frequency. Our approach is valid for arbitrary dispersion profiles, including resonant regimes where Taylor expansions of the dispersion relation fail and does not rely on frame hopping, allowing for multiple-velocity systems and arbitrarily accelerated interfaces. This opens the door to realistic implementations, particularly in upcoming experiments involving ENZ materials in the optical regime.
II Frequency Transitions
We study the problem of electromagnetic scattering at a modulation interface moving with uniform velocity $v_{\text{m}}$ along the $z$-axis and separating two isotropic, linear and passive but dispersive media [^1] For simplicity, the media are also considered to be lossless. However, loss may be straightforwardly added to the dispersive responses.. Both media are characterized by frequency-dependent refractive indices, $n_{1,2}{\left[\omega\right]}$, which satisfy the dispersion relation in their related medium: $\beta_{1,2}^{2}{\left[\omega\right]}=n_{1,2}^{2}{\left[\omega\right]}\omega^{2}/c^{2}$. An incident pulse in the first medium impinges on the moving interface and generates reflected and transmitted waves whose frequencies differ from the incident frequency due to the combined effects of Doppler shift and material dispersion.
The spectral components of the scattered waves are obtained by imposing phase matching [Deck-Leger2024_Dispersion_CONF] between the scattered and the incident waves along the trajectory of the interface. Writing the fields with the time-harmonic convention $\exp\left(i\left(\beta z-\omega t\right)\right)$, phase continuity along $z=v_{\text{m}}t$ yields
$\beta_{1}{\left[\omega_{1}\right]}z-\omega_{1}t$ $=\left.\beta_{1}{\left[\omega_{\text{i}}\right]}z-\omega_{\text{i}}t\right|_{z=v_{\text{m}}t}\,$, (1a)
$\beta_{2}{\left[\omega_{2}\right]}z-\omega_{2}t$ $=\left.\beta_{1}{\left[\omega_{\text{i}}\right]}z-\omega_{\text{i}}t\right|_{z=v_{\text{m}}t}\,$, (1b)
where $\omega_{\text{i}}$, $\omega_{1}$ and $\omega_{2}$ are the incident, reflected and transmitted frequencies, respectively. Inserting the dispersion relation, the trajectory parametrization and eliminating the explicit time-dependence in Eqs. (1), we obtain the following characteristic equations for the frequency transitions as a function of the incident frequency:
$\left(1+n_{1}{\left[\omega_{1}\right]}\frac{v_{\text{m}}}{c}\right)\omega_{1}$ $=\left(1-n_{1}{\left[\omega_{\text{i}}\right]}\frac{v_{\text{m}}}{c}\right)\omega_{\text{i}}\,$, (2a)
$\left(1-n_{2}{\left[\omega_{2}\right]}\frac{v_{\text{m}}}{c}\right)\omega_{2}$ $=\left(1-n_{1}{\left[\omega_{\text{i}}\right]}\frac{v_{\text{m}}}{c}\right)\omega_{\text{i}}\,.$ (2b)
In the nondispersive limit, i.e., $n_{1,2}{\left[\omega\right]}=n_{1,2}$, these equations reduce to the standard reflected and transmitted Doppler shifts at a constant moving interface [Caloz2019b_ST_Metamaterials_PUB]. However, in the dispersive case, Eqs. (2) can admit multiple solutions for each medium, each corresponding to a distinct scattered mode. Therefore, we denote the frequency of the $j$-th scattered wave in medium $i$ by $\omega_{i}^{\pm(j)}$, where $\pm$ indicates forward (++) or backward (-) propagation and $j=1,2,\dots$ labels distinct scattered waves.
[…]
Table 1: Group-velocity restriction and media passivity conditions for physical frequency transitions.
Medium Forward (++) Backward (-)
[…]
III Application to Drude Media
III.1 Specific Drude Solutions
Let us apply the general methodology of the previous section to the lossless Drude model, in which the refractive index profiles of the two media are given by
$n_{1}{\left[\omega\right]}$ $=\sqrt{n_{\infty,1}^{2}-\frac{\omega_{\text{p},1}^{2}}{\omega^{2}}}\,$, (3a)
$n_{2}{\left[\omega\right]}$ $=\sqrt{n_{\infty,2}^{2}-\frac{\omega_{\text{p},2}^{2}}{\omega^{2}}}\,$, (3b)
where $n_{\infty,1,2}$ are the high-frequency refractive indices and $\omega_{\text{p},1,2}$ are the plasma frequencies. We consider the general scenario in which both the high-frequency refractive index and the plasma frequency may be simultaneously modulated. Inserting Eqs. (3) into Eqs. (2) yields two second-order polynomials in $\omega_{1,2}$, which can be easily solved as
$\omega_{1}=\frac{1}{1-n_{\infty,1}^{2}v_{\text{m}}^{2}/c^{2}}\bigg(\left(1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c\right)\omega_{\text{i}}$
$\hskip 113.81102pt\pm v_{\text{m}}/c\sqrt{n_{\infty,1}^{2}\left(1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c\right)^{2}\omega_{\text{i}}^{2}-\left(1-n_{\infty,1}^{2}v_{\text{m}}^{2}/c^{2}\right)\omega_{\text{p},1}^{2}}\bigg)\,$, (4a)
$\omega_{2}=\frac{1}{1-n_{\infty,2}^{2}v_{\text{m}}^{2}/c^{2}}\bigg(\left(1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c\right)\omega_{\text{i}}$
$\hskip 113.81102pt\pm v_{\text{m}}/c\sqrt{n_{\infty,2}^{2}\left(1-n_{1}{\left[\omega_{\text{i}}\right]}v_{\text{m}}/c\right)^{2}\omega_{\text{i}}^{2}-\left(1-n_{\infty,2}^{2}v_{\text{m}}^{2}/c^{2}\right)\omega_{\text{p},2}^{2}}\bigg)\,.$ (4b)
Equations (4) always yield two candidates for the reflected solutions and two candidates for the transmitted solutions, as expected from the second-order characteristic polynomial [Eq. (2)]. To sort out the physical solutions from these four candidate solutions, we apply the conditions summarized in Table 1. Depending on the incident frequency ($\omega_{\text{i}}$), the modulation velocity ($v_{\text{m}}$) and dispersion parameters ($n_{\infty,1,2}$ and $\omega_{\text{p},1,2}$), the number and type of scattered waves vary. We refer to these distinct sets of allowed scattered waves as dispersion-mediated space-time states. Figure 1 illustrates these states for the lossless Drude model [Eq. (3)], with Fig. 1a demonstrating the different scattering regimes as a function of the modulation velocity and the incident frequency and Fig. 1b providing a space-index perspective and related spectral transition diagrams for each space-time state. Overall, ten different regimes, labeled A through J, are identified.
Figure 1: Dispersion-mediated space-time states for the lossless Drude model [Eq. (3)] with $n_{\infty,1}=1$, $\omega_{\text{p},1}=5$, $n_{\infty,2}=1.5$ and $\omega_{\text{p},2}=10$, computed according to the conditions in Tab. 1. (a) Scattering regimes of the space-time states as a function of the modulation velocity, $v_{\text{m}}$, and the incident frequency, $\omega_{\text{i}}$. The dashed lines indicate the nondispersive velocity limits. (b) Space-index perspective for each space-time state with corresponding spectral transition diagrams, where some of the solutions of Eq. (4) are crossed out since they do not satisfy the conditions in Tab 1.
III.2 Conventional Modes
First, we identify the regions corresponding to known nondispersive scattering regimes [DeKinder2025_Scat_Chirp_ASTEM], which are also recovered at high frequencies where dispersion becomes negligible. Region C in Fig. 1 corresponds to the conventional subluminal regime, where an incident wave ($\psi_{\text{i}}$) generates a backward reflected ($\psi_{1}^{-(1)}$) and a forward transmitted wave ($\psi_{2}^{+(1)}$). No backward transmitted wave ($\psi_{2}^{-(1)}$) appears because such a solution would violate the group-velocity restriction: the corresponding wave would propagate slower (larger in absolute value) than the interface and would therefore be overtaken and annihilated, rendering the solution unphysical. In the spectral transition diagram, this condition is visualized by comparing the slope of the dispersion curve at the intersection point, which corresponds to the group velocity, with the slope of the transition line, which corresponds to the modulation velocity. Region B corresponds to a contradirectional interluminal regime, where an incident wave produces reflected, later-backward and transmitted waves. Only in this regime are all the solutions in Eq. (4) physically valid, as they all satisfy both the group-velocity and passivity conditions. Region D represents the codirectional interluminal regime, in which the incident wave produces only a reflected wave. In contrast to the nondispersive case, this regime can also occur for contradirectional interfaces ($v_{\text{m}}<0$). This regime can be seen as total internal reflection introduced in [Agrawal2015_Temporal_Analog_USTEM_PUB]. Finally, regions A and E correspond to superluminal regimes. In the codirectional superluminal regime (region E), there is no scattering as the interface moves too fast for the incident wave to catch up. In the contradirectional superluminal regime (region A), an incident wave generates a later-forward and a later-backward wave. There is no forward scattered wave in the first medium ($\psi_{1}^{-(1)}$) due to the group-velocity restriction. An important difference with the nondispersive limit is that the boundaries between these regimes become frequency dependent due to group-velocity dispersion. In nondispersive systems, the regime boundaries depend only on the modulation velocity. Dispersion shifts and reshapes these boundaries and even extends some of the regimes beyond their nondispersive counterparts, as for the codirectional interluminal region D.
III.3 Unconventional Modes
In addition to the conventional, nondispersive-like regimes A–E, we uncover a set of purely dispersive space-time modes, which have no analog in nondispersive systems. These new modes correspond to regions F through J in Fig. 1. For regions F, G and H, the incident frequency is relatively close to the plasma frequency of the first medium, while for regions I and J, $\omega_{\text{i}}$ may lie relatively far from $\omega_{\text{p},1}$.
Regions F, G and H occur for a contradirectional moving interface ($v_{\text{m}}<0$). Starting from region H, scattering consists of a reflected ($\psi_{1}^{-(1)}$) and a later-backward ($\psi_{2}^{-(1)}$) wave. No later-forward wave ($\psi_{2}^{+(1)}$) is present because the transition line does not intersect with the dispersion curve of the second medium with a positive slope, so no physical forward mode in medium 2 is allowed. As the magnitude of the velocity is decreased—moving from region H to region G—the transition line couples with a negative slope intersection point of the dispersion diagram of the second medium. This give rise to a second later-backward wave ($\psi_{2}^{-(2)}$), in addition to the reflected and first later-backward wave. Notably, the group velocity of this new wave is very small, as indicated by the nearly horizontal slope of the dispersion curve at the intersection point. This suggests that the wave remains localized near its scattering position, effectively forming a quasi-stationary mode bound to the vicinity of its scattering position. Further decreasing the velocity into region F, the reflected wave is eliminated by the group-velocity restriction, as its trajectory is now overtaken by the interface.
Regions I and J are particularly interesting because they can exist even when the incident frequency is not close to the plasma frequency of the first medium. These regions host a forward wave in the first medium ($\psi_{1}^{+(1)}$), as the transition line intersects the dispersion curve of the first medium with a positive slope. This necessarily implies that the group velocity of this mode is smaller than the modulation velocity; otherwise, it would be excluded by the group-velocity restriction. In region I, the transmitted waves turn out to have complex frequencies and wavenumbers that violate the passivity constraints, and are therefore ruled out, leaving only reflected and forward waves in the first medium. Overall, the dispersion-mediated space-time modes in Fig. 1 emerge because the dispersion curves allow multiple intersections with the transition line. Each additional intersection corresponds to a new scattered mode, provided it satisfies group-velocity and passivity constraints. These additional modes do not appear in the nondispersive limit, where the dispersion curves are linear and intersect the transition line at most once per medium and direction.
IV Application to Lorentz Media
The methodology presented in the second section is applicable to any dispersive model, such as e.g. the Lorentz dispersive model, where $n{\left[\omega\right]}=\sqrt{n_{\infty}^{2}+\omega_{\text{p}}^{2}/\left(\omega_{0}^{2}-i\gamma\omega-\omega^{2}\right)}$, where $\omega_{\text{p}}$, $\omega_{0}$ and $\gamma$ are the plasma frequency, resonance frequency and damping factor, respectively. In this case, the characteristic equations [Eq. (2)] become fourth-order polynomials, potentially yielding up to four distinct solutions and thus a considerably more complex and richer scattering landscape. Figure 2 shows the different regions for a Lorentz temporal dispersion example, allowing up to $21$ regimes. As before, some nondispersive scattering regimes reappear, but are enriched by dispersion-mediated space-time states, which gradually reduce to the nondispersive behavior at high frequencies. These Lorentzian space-time states further demonstrate that dispersion not only perturbs nondispersive scattering but can fundamentally restructure the scattering landscape, creating new classes of modes and transitions that disappear in the nondispersive limit.
Figure 2: Dispersion-mediated space-time states for the lossless Lorentz dispersion model with parameters $n_{\infty,1}=1,\omega_{\text{p},1}=4,\omega_{0,1}=3,n_{\infty,2}=1.5,\omega_{\text{p},2}=8$ and $\omega_{0,2}=3$, computed according to the conditions in Tab. 1. The dashed lines indicate the nondispersive velocity limits.
V Conclusions
We have presented a general methodology for determining the spectral content of waves scattered at a uniformly moving interface between dispersive media. By combining phase-matching conditions with two conditions—group-velocity restriction and passivity—we have identified the physically allowed frequency transitions among the multiple solutions of the characteristic equations. A key feature of this methodology is that it is valid for arbitrary dispersion profiles, including regions near resonances, without relying on a Taylor expansion of the dispersion relation. It also does not require frame hopping, which makes it suitable for systems with multiple velocities or accelerated interfaces. Applying this framework to the lossless Drude model, we have showed that, in addition to the conventional nondispersive modes, a variety of dispersion-mediated space-time modes emerge. These modes have no counterpart in nondispersive scattering problems and arise from additional intersections between the dispersion curves and the phase-matching transition line. Similar behavior is observed in Lorentz media, where the higher-order characteristic equations and resonant dispersion yield an even richer set of space-time states, with up to $21$ distinct scattering regimes. These results may provide a foundational framework for analyzing realistic space-time varying systems, especially those relying on strongly dispersive materials such as epsilon-near-zero platforms in the optical regime [Boyd2020_Time_Refr_ENZ_TEM_PUB, Segev2023_Single_Cycle_TEM_PUB, Kinsey2025_ST_Knife_TEM_PUB]. By explicitly accounting for dispersion and its role in creating new frequency-transition branches, our approach offers a systematic route to dispersion-engineer space-time metamaterials.
Future works includes determining the scattering coefficients associated with each space-time state, thereby providing a complete solution to the electromagnetic scattering problem at a constant moving modulation interface in dispersive media. This task is nontrivial because it involves two distinct notions of time: an “interface time,” associated with the moving boundary, and a “memory time,” associated with the dispersive response of the medium. Dispersion is conventionally treated in the Fourier domain, whereas moving-boundary scattering problems are typically formulated in the time domain, making their consistent combination challenging. Other promising avenues for future research include extending the presented methodology to accelerated interfaces or multiple-velocity systems [Deck-Leger2019_Uni_Vel_PUB, Bahrami2025_Wedges_USTEM_PUB, Luo2025_ST_Super_Focu_USTEM_PUB].
Acknowledgments
K.D.K. is supported by the Research Foundation – Flanders (FWO) doctoral fellowship 1174526N.